Sol Summary – March 10th

Crew 293 Sol Summary Report 10-03-2024

Sol: 21

Summary Title: A Sunday Quest

Author’s name: Marie Delaroche, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

Today was supposed to be a rest day… but was it really? Many factors would indicate that it was in fact not. An hour of sleep was lost tonight, because apparently, Mars also has a daylight savings time change. This morning, after an hour-long workout, experiments had to be run in the Science Dome, and the Crew Astronomer had much troubleshooting to do in the Observatory.

In the afternoon, we knew that the HSO and Crew Engineer had something special planned for us. What we didn’t know, is that is turned out to be a station-wide treasure hunt! For the past week, they had been meticulously preparing this Sunday adventure for the rest of the Crew. We had so much fun solving riddles and passing trials, that we didn’t see time fly by.

We are looking forward to a restful evening and a good night’s sleep.

Look Ahead Plan:

The first EVA at Kissing Camel Ridge is planned tomorrow, to choose checkpoints and perform photogrammetry of the site.

Anomalies in work: Solar Observatory dome, EVA suits 7 and 9

Weather: Partly cloudy, Low 0°C, High 15°C

Crew Physical Status: Nominal


Reports to be filed: Sol Summary, Journalist’s Report, GreenHab Report, Operations Report, Photos (6-8 pics), EVA Request, Astronomy Report

Support Requested: NA

GreenHab Report – March 9th

Crew 293 GreenHab Report 09-03-2024

GreenHab Officer: Mathurin FRANCK

Environmental control: (heater, fan, door open?) Heater/fan turning on automatically

Average temperatures: between 70°F- 100°F

Hours of supplemental light: 10pm-2am

Daily water usage for crops: 8G

Daily water usage for research and/or other purposes: None

Water in Blue Tank (200 gallon capacity): 164 gallons

Time(s) of watering for crops: 10:00 am 8:00pm

Changes to crops: /

Narrative: /

Harvest: (include which crop and mass in grams): /

Support/supplies needed: /

Operations Report – March 9th

Crew 293 Operations Report 09-03-2024

SOL: 20

Name of person filing report: Leo Tokaryev

Non-nominal systems:

Notes on non-nominal systems:


Spirit rover used:no


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:yes


Beginning charge:100

Ending charge:44

Currently charging:yes

Curiosity rover used: yes

Hours: 270.8h

Beginning charge:100

Ending charge:43

Currently charging:yes

Perseverance rover used: no


Beginning charge:

Ending charge:

Currently charging:

General notes on rovers:

Summary of Hab operations: put notes here

WATER USE: 12.5 gallons

Water (static tank): 195.5 gallons

Static tank pipe heater (on or off):On

Static tank heater (On or off)On

Toilet tank emptied: no

Summary of internet: none

Summary of suits and radios:

Summary of GreenHab operations: put notes here

WATER USE: 8 gallons

Heater: On

Supplemental light: 22h-2h automatic

Harvest: none

Summary of ScienceDome operations: none

Dual split: Heat or AC, On

Summary of RAM operations: none

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support: we have left 2 trash bags in the rear airlock, could you please take it out? Thanks!

Journalist Report – March 9th

SOL 20: Back to the Canyon

“I’ve crossed these sands many times,” said one of the camel drivers one night. “But the desert is so huge, and the horizons so distant, that they make a person feel small, and as if he should remain silent.” – The Alchemist, Paulo Coelho

This morning, Crew 293’s last EVA to Candor Chasma took place! Léa and I had to find 12 checkpoints scattered all along the canyon. We prepared the expedition yesterday thanks to a 2D map and a topographic map. Mathurin was with us during the EVA: his goal was to count our steps, to take notes about our timing and performance. As we were getting closer to the site with the rovers, Mathurin had to stop his rover sooner than Léa and I, because his battery had drained faster than ours, and he would not have been able to come back to the station if he had continued. Léa and I continued driving Opportunity, to get as close as possible to Candor Chasma, and to limit our walking time to the canyon. Of course, we kept Mathurin in our line of sight at all times. Even though we found almost all of the checkpoints in less than an hour and a half, the 3D team was more efficient yesterday… But we are very proud to have managed to get this far, and we performed better than last week at North Ridge! Plus, going to Candor Chasma means enjoying its exceptional geology. Upon arriving at the end of the delimited search area, we were amazed to notice how far the canyon extended. The large rock tours surrounding us made us feel tiny compared to the Martian terrain’s wideness!

Back from the EVA, we were very tired. A 4-hour hike under the Martian Sun, while climbing on rock walls, is exhausting! We were lucky to just have to sit at the table to eat a plateful of good mashed potatoes from MELiSSA that was waiting for us! During lunch, we recounted our EVA to the rest of the Crew, and Crew members who stayed at the station brought us up to speed as well. Yves, who was our HabCom, cooked a brownie for dessert, Leo took care of the GreenHab, and Marie and Lise confronted both the software and hardware problems of the EchoFinder experiment. Everyone was productive during the absence of the EVA Crew!

During the afternoon, after some Crew members rested for a while, everyone followed its planned activities. Léa made impressive solar observations for her astronomy project. Yves and Marie tried once again to make EchoFinder’s equipment work for another session, which wasn’t very successful…

For the past few days, Lise and Leo have been preparing in secret an activity for tomorrow, that they finished today. We can’t wait to see what it is, and to enjoy tomorrow to rest. We need to be in good shape for our last week on Mars!

Astronomy Report – March 9th

[category  astronomy-report]

Astronomy Report

Name: Léa Bourgély

Crew: 293

Date: 09/03/2024


Telescope Requested: Solar Observatory

Objects to be Imaged this Evening: None

Images submitted with this report: Sun prominence

Additional Notes:

1. The lower part of the shutter is still stuck, but we’ve got some improvements. Peter shared with us new guidelines for tomorrow, we will apply them during the morning. We are making progress.

2. About imaging with the solar telescope : here is a nice picture of a prominence, and it disappeared some minutes later. I’ve been trying to get good images of sunspots (like the ones that are present in the guide book) but there is no "GAMMA" to tune on SharpCap, and I am not able to tune really well the telescope with only exposure and gain. @peter, any piece of advice on that ? Also, I think tomorrow’s conditions for imaging are better than today’s. I am also wondering how to tune the telescope with the 2 big black knobs ? I’ve been trying them but I can’t see the sun’s surface like the one present in the guide books. Thanks for your help !!

EVA Report – March 9th

Crew 293 EVA Report 09-03-2024

EVA 14

Author: Léa Bourgély

Purpose of EVA: 2D Photogrammetry EVA at Candor Chasma

Start time: 09:00

End time:13:00

Suits and radio check:

Suit 1: 13.35V 2.6LPM

Suit 2: 13.3V 2.4LPM

Suit 3: 13.35V 2.6LPM

Radio 3: 8.09V

Radio 6: 8.09V

Radio 1: 7.99V

Radio 7 (HabCom): 8.09V


08:59 Depressurization begins

09:04 EVA begins

Curiosity: 270.8h, 100%

Opportunity: 166.4 100%

09:08 Leaving campus

09:10 Arrived at the atmospheric site

09:33 Data retrieved (not the field mill), batteries changed. Heading back to the rovers

09:35 Leaving atmospheric site. Switching radio to channel 1.

09:47 Leaving the rovers.

10:00 Arriving at the entrance of Candor Chasma. Contacting every 15min to make sure everything is alright.

10:12 Arriving in the canyon

10:30 almost at the starting point

10:36 Reached the starting point

12:04 Going back to the rovers

12:37 Back to the rovers.

12:39 Leaving site

12:47 Arriving on campus

Curiosity 43%

Opportunity 44%

12:50 Pressurization begins

12:55 End of EVA

Suits checkup: (END)
Suit 1: 12.50V

Suit 2: 12.50V

Suit 3: 12.69V

Destination: Candor Chasma

Coordinates (use UTM WGS 84): 520500, 4251000

Participants: Léa(EVA Leader), Erin and Mathurin

Road(s) and routes per MDRS Map: Cow Dung Road, North direction, then Galileo Road to Compass Rock

Mode of travel: Rovers and walking

Sol Summary – March 9th

Crew 293 Sol Summary Report 09-03-2024

Sol: 20

Summary Title: Back to the Canyon

Author’s name: Marie Delaroche, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

Although it is never not special to wake up to the Martian hills drenched in rising sunlight, it is also increasingly difficult to get out of bed at 6:45… This morning’s challenge was the final EVA to Candor Chasma, this time for the Crew Journalist and Crew Astronomer to reach the checkpoints with the sole help of a 2D map. While they did come back with wide smiles, the exertion finally caught up, and when I came back from the Science Dome after an EchoFinder session, the Hab was doused in silence: the Crew was resting, a rest that felt as though it had been long awaited.

Some good news: AMI is finally fully operational. The interface now has the possibility to record and store the actions performed by the Crew in power management. This allows us to run the interface and manage the station’s resources without the additional task of manually recording major configuration changes. Unfortunately, regarding EchoFinder, we are having more and more issues with both the software and the hardware. We are awaiting Earth-based help to try and solve these issues, although we have already been able to gather significant amounts of data.

Finally, despite issues with the shutter system of the observatory, the Crew Astronomer was able to photograph solar prominences and more sunspots for her study.

Look Ahead Plan:

Sol 21 has been declared a rest day. For our final week on Mars, three EVAs are planned at Kissing Camel Ridge, the final Photogrammetry site (Sol 22, 24, 25). We aim to dismantle and retrieve all atmospheric instruments on Sol 26.

Anomalies in work: Solar Observatory dome

Weather: Sunny, low -5°C, high 15°C

Crew Physical Status: Nominal

EVA: Exploration EVA to Candor Chasma

Reports to be filed: Sol Summary, Journalist’s Report, GreenHab Report, Operations Report, Photos (6-8 pics), EVA Report, Astronomy Report

Support Requested: NA

Journalist Report – March 11th

SOL 22: Good Vibes Only!

“Don’t forget that everything you deal with is only one thing and nothing else. And don’t forget the language of omens. And, above all, don’t forget to follow your Personal Legend through to its conclusion.” – The Alchemist, Paulo Coelho

To start the week on a good note, Yves and Mathurin initiated the last step of the Photogrammetry experiment. During their EVA, they went to Kissing Camel Ridge in order to take pictures of the area and to generate 2D and 3D maps of it. Twice we explored the northern part of the station’s surroundings, but this time, we turned right after leaving campus, taking the southbound road! It enabled us to discover new Martian landscapes… During this EVA, Yves and Mathurin needed to be with another Crew member. This time it was Lise, instead of Leo the two previous weeks. Lise, who had loved searching for the checkpoints in the previous destinations, was on the other side of the experiment this time, and she loved it all the same! Her goal, with Yves, was to find where to put the checkpoints while Mathurin was taking the hundreds of drone pictures. During the week, there will be new exploration teams with the 2D or 3D map. This is the last iteration of an experiment we all very much enjoyed participating in!

Meanwhile, Léa and Marie wrote some new outreach articles for our website, and every Crew member who wasn’t out on EVA was busy accomplishing their planned tasks, from experimental data handling to cognitive assessments in the Hab or GreenHab.

The MELiSSA activity, prompting us to cook meals with fresh vegetables and not only dehydrated food, continues to delight the Crew! Today for example, Léa and I cooked a very tasty carrot cake for lunch. Crew members who were out on EVA enjoyed this great meal after being out for a few hours. Then, in the early afternoon, Lise and Mathurin tried to solve problems they had in their code for retrieving data from our smart watches and impedance meter. They put in common everything they had done on their own. At the end of the afternoon, we all enjoyed a virtual reality experience: in this personalized video for each of us, we were surprised to see Arnaud Prost, a French astronaut, wishing us good luck and giving us advice for the end of our mission! This video cheered everyone up, and gave us that final bit of motivation we needed, to end all of our studies and experiments in a grand finale!

GreenHab Report – March 11th

Crew 293 GreenHab Report 11-03-2024

GreenHab Officer: Mathurin FRANCK

Environmental control: (heater, fan, door open?) Heater/fan turning on automatically

Average temperatures: between 70°F- 85°F

Hours of supplemental light: 10pm-2am

Daily water usage for crops: 8G

Daily water usage for research and/or other purposes: None

Water in Blue Tank (200 gallon capacity): 148 gallons

Time(s) of watering for crops: 10:00 am 5:00pm

Changes to crops: /

Narrative: /

Harvest: (include which crop and mass in grams): /

Support/supplies needed: /

Copyright © The Mars Society. All rights reserved. | Main Site